# Lane Graph as Path: Continuity-preserving Path-wise Modeling for Online Lane Graph Construction

### v2

@@ Abstract
Online lane graph construction is a promising but challenging task in autonomous driving. Previous methods usually model the lane graph at the pixel or piece level, and recover the lane graph by pixel-wise or piece-wise connection, which breaks down the continuity of the lane. Human drivers focus on and drive along the continuous and complete paths instead of considering lane pieces. Autonomous vehicles also require path-specific guidance from lane graph for trajectory planning. We argue that the path, which indicates the traffic flow, is the primitive of the lane graph. Motivated by this, we propose to model the lane graph in a novel path-wise manner, which well preserves the continuity of the lane and encodes traffic information for planning. We present a path-based online lane graph construction method, termed LaneGAP, which end-to-end learns the path and recovers the lane graph via a Path2Graph algorithm. We qualitatively and quantitatively demonstrate the superiority of LaneGAP over conventional pixel-based and piece-based methods on challenging nuScenes and Argoverse2 datasets. Abundant visualizations are available in the supplementary material, which show LaneGAP can cope with diverse traffic conditions. Code and models will be released for facilitating future research.
footnotetext: ${}^{\star}$ Equal contribution; ${}^{\diamond}$ Interns of Horizon Robotics when doing this work; ${}^{\boxtimes}$ Corresponding author: xgwang@hust.edu.cn
@@ 1 Introduction
An intuitive solution is to model the lane graph in a pixel-wise manner and adopt a segmentation-then-vectorization paradigm. For example, HDMapNet [20] predicts a segmentation map and a direction map on dense bird’s-eye-view (BEV) features. It then extracts the skeleton from the coarse segmentation map with a morphological thinning algorithm and extracts the graph topology by greedily tracing the single-pixel-width skeleton with the predicted direction map. Pixel-wise modeling incurs heuristic and time-consuming post-processing and often fails in complicated topology (see Fig. 2 row 3).
Lane graph is also modeled in a piece-wise manner [6], which splits the lane graph into lane pieces at junction points (i.e., merging points and fork points) and predicts an inter-piece connectivity matrix. Based on the connectivity, pieces are linked and merged into the lane graph. However, piece-wise modeling breaks down the continuity of the lane. Aligning the pieces is a challenging problem, especially at complicated road intersections. And piece-wise modeling often breaks down the lane graph into fragmented short pieces which are hard for a neural network to learn to perceive (e.g., $V_{1}^{\text{piece}}$ in Fig. 1 (b)).
We argue that the path is the primitive of the lane graph. Human drivers focus on and drive along the continuous and complete paths instead of considering lane pieces. Autonomous vehicles also require path-specific guidance from lane graph for trajectory planning. Continuous paths play an important role in indicating the driving flow, while pixel-wise and piece-wise modeling often fail to merge pixels and pieces into continuous paths (see Fig. 2).
@@
Figure 1: Modeling comparison. (a) Pixel-wise modeling [20] utilizes a predefined Graph2Pixel algorithm to rasterize the lane graph into a segmentation map and a direction map on dense BEV pixels, and heuristic Pixel2Graph post-processing is needed to recover the lane graph from the predicted segmentation map $V_{\text{pixel}}$ and direction map $D_{\text{pixel}}$ (direction map is not drawn here for simplicity). (b) Piece-wise modeling [6] utilizes a predefined Graph2Piece algorithm to split the lane graph into a set of pieces and the connectivity matrix among pieces, and then it merges the predicted pieces $\mathcal{V}_{\text{piece}}$ to the graph with the Piece2Graph algorithm based on predicted connectivity $E_{\text{piece}}$. (c) The proposed path-wise modeling translates the lane graph into continuous and complete paths with a simple predefined Graph2Path algorithm to traverse the graph. We perform path detection and adopt a simple Path2Graph algorithm to recover the lane graph.
We compare LaneGAP with the pixel-wise modeling method HDMapNet [20] and piece-wise modeling methods STSU [6], MapTR [23] under strictly fair conditions (encoder, model size, training schedule, etc.) on the challenging nuScenes [5] and Argoverse2 [41] datasets, which covers diverse graph topology and traffic conditions. With only camera input, LaneGAP achieves the best graph construction quality both quantitatively (see Tab.1) and qualitatively (see Fig. 2), while running at the fastest inference speed. We further push forward the performance of lane graph construction, by introducing the multi-modality input. We believe modeling the lane graph at the path level is reasonable and promising. We hope LaneGAP can serve as a fundamental module of the self-driving system and boost the development of downstream motion planning.
@@
  • We qualitatively and quantitatively demonstrate the superiority of LaneGAP over pixel-based and piece-based methods. LaneGAP can cope with diverse traffic conditions, especially for road intersections with complicated lane topology.
Figure 2: Qualitative comparison of pixel-wise HDMapNet, piece-wise MapTR, and path-wise LaneGAP on the nuScenes and Argoverse2 val splits. Given the ${}^{\circ}$ horizontal FOV images as input, the proposed path-wise modeling well preserves the continuity across different scenarios, especially for complicated lane topology with more than 4 junction points. The top 2 rows are from the nuScenes val split, and the bottom 2 rows are from the Argoverse2 val split. More visualizations are available in the supplementary material.
@@ 2 Related Work
Lane detection.
Lane detection only considers predicting and evaluating the lane divider lines without spatial relation (merging and fork). Since most lane detection datasets only provide front-view images, previous lane detection methods [38, 40, 15, 26, 16, 25] were stuck in predicting lines with a small curvature in a limited horizontal FOV. BezierLaneNet [14] uses a fully convolutional network to predict Bezier lanes defined with 4 Bezier control points. PersFormer [9] proposes a Transformer-based architecture for spatial transformation and unifies 2D and 3D lane detection.
Online HD map construction. Online HD map construction can be seen as an advanced setting of lane detection, consisting of lines and polygons with various semantics in the local ${}^{\circ}$ FOV perception range of ego-vehicle. With advanced 2D-to-BEV modules [32], previous online HD map construction methods cast it into semantic segmentation task on the transformed BEV features [29, 45, 21, 30, 28, 31]. Building vectorized semantic HD map online achieves increasing interests nowadays [20, 23, 27, 37, 35, 12], HDMapNet [20] follows a segmentation-then-vectorization paradigm. To achieve end-to-end learning [8, 47, 13], VectorMapNet [27] adopts a coarse-to-fine two-stage pipeline for vectorized HD map learning. MapTR [23] proposes unified permutation-equivalent modeling to exploit the undirected nature of semantic HD map and designs a parallel end-to-end framework. BeMapNet [35] and PivotNet [12] propose a Bezier-based representation and pivot-based representation for modeling the map geometry. While the above works focus on the map elements without physical directions (i.e., lane divider, pedstrian crossing and road boundary) and lane graph topology, we aims to fill the gap in this work.
@@
Lane graph construction. Lane graph is traditionally constructed with an offline pipeline [42, 18, 3]. [18] proposes a mutli-step training pipeline to construct the lane graph of aerial images based on pixel-wise modeling. [3] proposes a bottom-up approach to aggregate multiple local aerial lane graphs into a global consistent graph. CenterlineDet [42] proposes a DETR-like decision-making transformer network to iteratively update the global lane graph with vehicle-mounted sensors. Recently, STSU [6] shift the offline lane graph construction to the online, ego-centric setting with vehicle-mounted sensors. It models the lane graph as a set of disjoint pieces split by junction points and a set of connections among those pieces. A DETR-like Transformer decoder is proposed to detect those Beizer lane pieces and a successive MLP head is used to predict inter-piece connectivity. Based on STSU, [7] designs a network and utilizes minimal circle extracted by time-consuming offline processing to further supervise the network to produce lane graph. Different from the above graph modelings, we regard the path as the primitive of the lane graph, and model the lane graph in a novel path-wise manner.
@@ 3 Method
Figure 3: Overview of LaneGAP. Based on the proposed path-wise modeling, LaneGAP first utilizes the Graph2Path algorithm to translate the graph into a set of paths. Then, an end-to-end network is designed to represent and learn the paths. Finally, a Path2Graph algorithm is used to translate the predicted paths into graph structure with negligible cost. Graph2Path and Path2Graph are deterministic algorithms and don’t require backpropagation for optimization.
In this section, we first describe how to translate the directed lane graph into a set of directed paths in Sec. 3.1. Then we introduce the online path detection framework in Sec. 3.2. And we describe how to translate paths back to the lane graph in Sec. 3.3. An overview of our method is exhibited in Fig. 3.
@@ 3.1 Graph2Path
Input: directed lane graph $G$
Output: directed paths $\mathcal{V}$
1 Initialization: $\mathcal{V}\leftarrow\emptyset$, root vertices set $\text{root vertices set }\Omega_{\text{root}}\leftarrow\emptyset$, leaf vertices set $\text{leaf vertices set }\Omega_{\text{leaf}}\leftarrow\emptyset$ for vertice in $G$.vertices do
[…]
3             $\Omega_{\text{root}}\leftarrow\Omega_{\text{root}}\cup\{vertice\}$
4       end if
[…]
6             $\Omega_{\text{leaf}}\leftarrow\Omega_{\text{leaf}}\cup\{vertice\}$
7       end if
8
9 end for
10for root vertice in $\Omega_{\text{root}}$  do
11       for leaf vertice in $\Omega_{\text{leaf}}$ do
12             $V^{\text{path}}=\textbf{findpath}(root\ vertice,leaf\ vertice,G)$ if  $V^{\text{path}}\text{ is not None}$ then
13                   $\mathcal{V}\leftarrow\mathcal{V}\cup\{V^{\text{path}}\}$
14             end if
15
16       end for
17
18 end for
Return: $\mathcal{V}$
Algorithm 1 Pseudo-code of Graph2Path.
@@ 13                   $\mathcal{V}\leftarrow\mathcal{V}\cup\{V^{\text{path}}\}$
Given the ground truth Lane graph $G$, which is typically a directed graph in the local map around the ego-vehicle, we first extract the root vertices $\Omega_{\text{root}}$ and leaf vertices $\Omega_{\text{leaf}}$ based on the indegree and outdegree of vertice. Then we pair the root vertices and leaf vertices (line 2 to 9 in Alg. 1.). For each vertice pair, a depth-first-search (DFS) algorithm is used to find the valid path $V^{\text{path}}$ from the root vertice to the leaf vertice (line 10 to 17 in Alg. 1.). Finally, we can translate the ground truth directed lane graph $G$ to a set of directed paths $\mathcal{V}_{\rm path}=\{V_{i}^{\rm path}\}_{i=1}^{M}$, where $M$ is the number of ground truth paths.
@@ 3.2 Path Representation and Learning
Inspired by the advanced set detection methods [8, 47], we propose an end-to-end network, LaneGAP, to predict all paths simultaneously in a single stage, as illustrated in Fig. 3. Our network consists of an encoder that encodes the features from the onboard sensor data, and a query-based Transformer decoder that performs set detection by decoding a set of paths $\hat{\mathcal{V}}_{\rm path}=\{\hat{V}_{i}^{\rm path}\}_{i=1}^{N}$ from the encoded features. To parameterize the path, we utilize two widely-used types of representations, Polyline [27, 23] and Bezier [6, 35], where Polyline offers high flexibility in describing the path, while Bezier provides a smoother representation.
@@
$\mathcal{L}_{\rm bipartite}(\hat{\mathcal{V}}_{\text{path}},\mathcal{V}_{\text{path}})=$ $\sum_{i=1}^{N}[\mathcal{L}_{\rm{Focal}}(\hat{p}_{{\hat{\sigma}}(i)},c_{i})+$ (3)
$\mathds{1}_{\{c_{i}\neq\varnothing\}}\mathcal{L}_{\rm path}(\hat{V}_{\hat{\sigma}(i)}^{\rm path},V_{i}^{\rm path})],$
@@
where $\hat{\sigma}$ is the optimal assignment between a set of predicted paths and a set of ground truth paths computed by the Hungarian algorithm, $c_{i}$ is the target class label, and $\mathcal{L}_{\rm{Focal}}(\hat{p}_{{\hat{\sigma}}(i)},c_{i})$ is the classification loss defined in [24]. A $L_{1}$ loss is utilized between the matched predicted path $\hat{V}_{\hat{\sigma}(i)}^{\rm path}$ and sampled ground truth path $V_{i}^{\rm path}$. To enhance the BEV features, we introduce an auxiliary BEV segmentation branch $\phi_{\rm seg}$ to predict the BEV segmentation mask, and the auxiliary BEV segmentation loss is defined as:
@@
3.3 Path2Graph
@@
1 Initialization: $G=\text{DirGraph}()$ for $V^{\text{path}}$ in $\mathcal{V}$ do
2       $V=\{v_{j}\}\leftarrow\textbf{discretize\_path\_to\_vertex\_sequence}(V^{\text{path}})$ for  $v_{j}$ in $V$ do
@@ 1 Initialization: $G=\text{DirGraph}()$ for $V^{\text{path}}$ in $\mathcal{V}$ do
4                   $G.\textbf{add\_vertice}(v_{j})$ break
@@
6            $G.\textbf{add\_vertice}(v_{j})$ $G.\textbf{add\_edge}(v_{j},v_{j+1})$
@@
$G=\textbf{merge\_vertices\_by\_overlap}(G)$ Return: $G$
@@
The predicted continuous paths encode sufficient traffic information and can be directly applied to downstream motion planning. To further recover the graph structure of lane topology and extract merging and fork information, we convert the predicted paths $\hat{\mathcal{V}}=\{\hat{V}_{i}^{\text{path}}\}$ into a directed lane graph $\hat{G}$ with a designed Path2Graph algorithm as shown in Alg. 2.
We discretize the path into point sequences $V=\{v_{j}\}$. The discretized points are regarded as vertices and the adjacent relation between successive points is regarded as edges of vertices. We add these vertices and edges to the directed graph (line 4 to 11 in Alg. 2). The vertices registered in the directed graph $G$ on one path may have spatial overlapping with vertices on other paths. To remove the redundancy of $G$, we merge the overlapped vertices into one vertex (line 13 in Alg. 2).

### v3

@@ Abstract
Online lane graph construction is a promising but challenging task in autonomous driving. Previous methods usually model the lane graph at the pixel or piece level, and recover the lane graph by pixel-wise or piece-wise connection, which breaks down the continuity of the lane and results in suboptimal performance. Human drivers focus on and drive along the continuous and complete paths instead of considering lane pieces. Autonomous vehicles also require path-specific guidance from lane graph for trajectory planning. We argue that the path, which indicates the traffic flow, is the primitive of the lane graph. Motivated by this, we propose to model the lane graph in a novel path-wise manner, which well preserves the continuity of the lane and encodes traffic information for planning. We present a path-based online lane graph construction method, termed LaneGAP, which end-to-end learns the path and recovers the lane graph via a Path2Graph algorithm. We qualitatively and quantitatively demonstrate the superior accuracy and efficiency of LaneGAP over conventional pixel-based and piece-based methods on the challenging nuScenes and Argoverse2 datasets under controllable and fair conditions. Compared to the recent state-of-the-art piece-wise method TopoNet on the OpenLane-V2 dataset, LaneGAP still outperforms by 1.6 mIoU, further validating the effectiveness of path-wise modeling. Abundant visualizations in the supplementary material show LaneGAP can cope with diverse traffic conditions. Code is released at https://github.com/hustvl/LaneGAP.
Keywords:
Autonomous Driving Online Lane Graph Construction Path-wise Modeling
[…]
@@ 1 Introduction
Figure 1: Modeling comparison. (a) Pixel-wise modeling [22] utilizes a predefined Graph2Pixel algorithm to rasterize the lane graph into a segmentation map and a direction map on dense BEV pixels, and heuristic Pixel2Graph post-processing is needed to recover the lane graph from the predicted segmentation map $V_{\text{pixel}}$ and direction map $D_{\text{pixel}}$ (direction map is not drawn here for simplicity). (b) Piece-wise modeling [6] utilizes a predefined Graph2Piece algorithm to split the lane graph into a set of pieces and the connectivity matrix among pieces, and then it merges the predicted pieces $\mathcal{V}_{\text{piece}}$ to the graph with the Piece2Graph algorithm based on predicted connectivity $E_{\text{piece}}$. (c) The proposed path-wise modeling translates the lane graph into complete paths with a predefined Graph2Path algorithm to traverse the graph. We perform path detection and adopt a Path2Graph algorithm to recover the lane graph.
@@
An intuitive solution is to model the lane graph in a pixel-wise manner and adopt a segmentation-then-vectorization paradigm. For example, HDMapNet [22] predicts a segmentation map and a direction map on dense bird’s-eye-view (BEV) features. It then extracts the skeleton from the coarse segmentation map with a morphological thinning algorithm and extracts the graph topology by greedily tracing the single-pixel-width skeleton with the predicted direction map. Pixel-wise modeling incurs heuristic and time-consuming post-processing and often fails in complicated topology (see Fig. 2 row 3, Tab. 10).
Lane graph is also modeled in a piece-wise manner [6], which splits the lane graph into lane pieces at junction points (i.e., merging points and fork points) and predicts an inter-piece connectivity matrix. Based on the connectivity, pieces are linked and merged into the lane graph. However, piece-wise modeling breaks down the continuity of the lane. At complicated road intersections, the fragmented short pieces are hard for a neural network to learn to detect (e.g., $V_{1}^{\text{piece}}$ in Fig. 1 (b)), let alone predict the connections.
We argue that the path is the primitive of the lane graph. Human drivers focus on and drive along the continuous and complete paths instead of considering lane pieces. Autonomous vehicles also require path-specific guidance from lane graph for trajectory planning [12]. Continuous paths play an important role in indicating the driving flow, while pixel-wise and piece-wise modeling often fail to merge pixels and pieces into continuous paths (see Fig. 2).
@@
Figure 2: Qualitative comparison of pixel-wise HDMapNet, piece-wise MapTR, and path-wise LaneGAP on the nuScenes and Argoverse2 val splits. The top 2 rows are from the nuScenes val split, and the bottom 2 rows are from the Argoverse2 val split. Different colors indicate different instances. More qualitative comparisons are available in the supplementary.
To fairly evaluate across different modelings and emphasize the unique properties of the lane graph (i.e., connection, direction, and split/merge around the junction points), we carefully design a graph-centric TOPO metric to evaluate both holistic topology and local junction topology, which is compatible with different modelings.
We compare LaneGAP with the pixel-wise modeling method HDMapNet [22] and piece-wise modeling methods STSU [6], MapTR [26] under strictly fair conditions (encoder, model size, training schedule, etc.) on the challenging nuScenes [5] and Argoverse2 [47] datasets, which covers diverse graph topology and traffic conditions. With only camera input, LaneGAP achieves the best graph construction quality both quantitatively (see Tab. 1) and qualitatively (see Fig. 2), while running at the fastest inference speed (2x faster than HDMapNet). We further validate the superiority of path-wise modeling on TopoNet [23] (1.6 higher mIoU), a state-of-the-art piece-wise modeling method on the OpenLane-V2 dataset [45]. The extensive experiments show that modeling the lane graph at the path level is reasonable and promising. We hope LaneGAP can serve as a fundamental module of the self-driving system and boost the development of downstream motion planning.
@@
  • We design a novel graph-centric metric to evaluate the unique properties of lane graph construction (i.e., connection, direction, and junction), which is compatible with different modelings.
  • We qualitatively and quantitatively demonstrate the superior accuracy and efficiency of LaneGAP over pixel-based and piece-based methods. LaneGAP can cope with diverse traffic conditions, especially for road intersections with complicated lane topology.
@@ 2 Related Work
Lane detection. Lane detection only considers predicting and evaluating the lane divider lines without spatial relation (merging and fork). Since most lane detection datasets only provide front-view images, previous lane detection methods [43, 46, 16, 29, 17, 28] were stuck in predicting lines with a small curvature in a limited horizontal FOV. BezierLaneNet [15] uses a fully convolutional network to predict Bezier lanes defined with 4 Bezier control points. PersFormer [9] proposes a Transformer-based architecture for spatial transformation and unifies 2D and 3D lane detection.
Online HD map construction. Online HD map construction can be seen as an advanced setting of lane detection, consisting of lines and polygons with various semantics in the local 360 FOV perception range of ego-vehicle. With advanced 2D-to-BEV modules [37], previous online HD map construction methods cast it into semantic segmentation task on the transformed BEV features [32, 53, 24, 33, 31, 36]. Building vectorized semantic HD map online achieves increasing interests nowadays [22, 26, 30, 42, 40, 13], HDMapNet [22] follows a segmentation-then-vectorization paradigm. To achieve end-to-end learning [8, 55, 14], VectorMapNet [30] adopts a two-stage pipeline for vectorized HD map learning. MapTR [26] proposes a unified permutation-equivalent modeling to exploit the undirected nature of semantic HD map and designs a parallel end-to-end framework. BeMapNet [40] and PivotNet [13] propose Bezier-based representation and pivot-based representation for modeling map geometry. While the above works focus on map elements without physical directions and lane graph topology, we aim to fill the gap in this work.
@@
Lane graph construction. The lane graph is traditionally constructed with an offline pipeline [49, 19, 3]. [19] proposes a multi-step training pipeline to construct the lane graph of aerial images based on pixel-wise modeling. [3] proposes a bottom-up approach to aggregate multiple local aerial lane graphs into a globally consistent graph. CenterlineDet [49] proposes a DETR-like decision-making transformer network to iteratively update the global lane graph with vehicle-mounted sensors. Recently, STSU [6] shifts the offline lane graph construction to the online, ego-centric setting with vehicle-mounted sensors. It models the lane graph as a set of disjoint pieces split by junction points and a set of connections among those pieces. Based on STSU, recent works [7, 23, 48] advance this piece-wise setting and push forward the performance. Different from the above graph modelings, we regard the path as the primitive of the lane graph, and model the lane graph in a novel path-wise manner.
@@ 3 Method
Figure 3: Overview of LaneGAP.
In this section, we first describe how to translate the directed lane graph into a set of directed paths in Sec. 3. Then we introduce the online path detection framework in Sec. 3.2. And we describe how to translate paths back to the lane graph in Sec. 3.3. An overview of our method is exhibited in Fig. 3.
@@ 3.1 Graph2Path
Given the ground truth Lane graph $G$, which is typically a directed graph in the local map around the ego-vehicle, we first extract the root vertices $\Omega_{\text{root}}$ and leaf vertices $\Omega_{\text{leaf}}$. Then we pair the root vertices and leaf vertices and then utilize depth-first-search (DFS) algorithm to find the valid path $V^{\text{path}}$. Finally, we can translate the ground truth directed lane graph $G$ to a set of directed paths $\mathcal{V}_{\rm path}=\{V_{i}^{\rm path}\}_{i=1}^{M}$, where $M$ is the number of ground truth paths.
@@ 3.2 Path Representation and Learning
Inspired by advanced set detection methods [8, 55], we propose an end-to-end network, LaneGAP, to predict all paths simultaneously in a single stage, as illustrated in Fig. 3. Our network consists of an encoder that encodes the features from the onboard sensor data, and a query-based Transformer decoder that performs set detection by decoding a set of paths $\hat{\mathcal{V}}_{\rm path}=\{\hat{V}_{i}^{\rm path}\}_{i=1}^{N}$ from the encoded features. To parameterize the path, we utilize two widely-used types of representations, Polyline [30, 26] and Bezier [6, 40], where Polyline offers high flexibility in describing the path, while Bezier provides a smoother representation.
@@
$\mathcal{L}_{\rm bipartite}(\hat{\mathcal{V}}_{\text{path}},\mathcal{V}_{\text{path}})=$ $\sum_{i=1}^{N}[\mathcal{L}_{\rm{Focal}}(\hat{p}_{{\hat{\sigma}}(i)},c_{i})+$ $\mathds{1}_{\{c_{i}\neq\varnothing\}}\mathcal{L}_{\rm path}(\hat{V}_{\hat{\sigma}(i)}^{\rm path},V_{i}^{\rm path})],$ (3)
@@
where $\hat{\sigma}$ is the optimal assignment between a set of predicted paths and a set of ground truth paths computed by the Hungarian algorithm, $c_{i}$ is the target class label, and $\mathcal{L}_{\rm{Focal}}(\hat{p}_{{\hat{\sigma}}(i)},c_{i})$ is the classification loss defined in [27]. $L_{1}$ loss is utilized between the matched predicted path $\hat{V}_{\hat{\sigma}(i)}^{\rm path}$ and sampled ground truth path $V_{i}^{\rm path}$. To enhance the BEV features, we introduce an auxiliary BEV segmentation branch $\phi_{\rm seg}$ to predict the BEV segmentation mask, and the auxiliary BEV segmentation loss is defined as:
@@
Input: directed lane graph $G$
Output: directed paths $\mathcal{V}$
1
2Initialize: $\mathcal{V}\leftarrow\emptyset$, root vertices set $\text{root vertices set }\Omega_{\text{root}}\leftarrow\emptyset$, leaf vertices set $\text{leaf vertices set }\Omega_{\text{leaf}}\leftarrow\emptyset$
3 for v in $G$.vertices do
[…]
5             $\Omega_{\text{root}}\leftarrow\Omega_{\text{root}}\cup\{v\}$
6
7       end if
[…]
9             $\Omega_{\text{leaf}}\leftarrow\Omega_{\text{leaf}}\cup\{v\}$
10
11       end if
12
13 end for
14for root v in $\Omega_{\text{root}}$  do
15       for leaf v in $\Omega_{\text{leaf}}$ do
16             $V^{\text{path}}=\textbf{findpath}(root\ v,leaf\ v,G)$
17             if  $V^{\text{path}}\text{ is not None}$ then
18                   $\mathcal{V}\leftarrow\mathcal{V}\cup\{V^{\text{path}}\}$
19
20             end if
21
22       end for
23
24 end for
25
Return: $\mathcal{V}$
Algorithm 1 Pseudo-code of Graph2Path.
@@ 18                   $\mathcal{V}\leftarrow\mathcal{V}\cup\{V^{\text{path}}\}$
1
2Initialization: $G=\text{DirGraph}()$
3 for $V^{\text{path}}$ in $\mathcal{V}$ do
4       $V=\{v_{j}\}\leftarrow\textbf{discret\_to\_vert\_seq}(V^{\text{path}})$
5       for  $v_{j}$ in $V$ do
@@ 4       $V=\{v_{j}\}\leftarrow\textbf{discret\_to\_vert\_seq}(V^{\text{path}})$
7                   $G.\textbf{add\_vertice}(v_{j})$
8                   break
9
@@
11            $G.\textbf{add\_vertice}(v_{j})$
12             $G.\textbf{add\_edge}(v_{j},v_{j+1})$
13
@@
[…]
Return: $G$
@@
3.3 Path2Graph
The predicted continuous paths encode sufficient traffic information and can be directly applied to downstream motion planning. To further recover the graph structure of lane topology and extract merging and fork information, we convert the predicted paths $\hat{\mathcal{V}}=\{\hat{V}_{i}^{\text{path}}\}$ into a directed lane graph $\hat{G}$ with a designed Path2Graph algorithm in Alg. 2.
We discretize the path into point sequences $V=\{v_{j}\}$. The discretized points are regarded as vertices and the adjacent relation between successive points is regarded as edges of vertices. We add these vertices and edges to the directed graph (line 4 to 11 in Alg. 2). The vertices registered in the directed graph $G$ on one path may have spatial overlapping with vertices on other paths. We merge the overlapped vertices into one vertex if their distance is below a threshold (line 13 in Alg. 2). The merged vertex inherits the predecessors and successors of the pair of overlapped vertices.